Jun is 5 years older than july. the sum of their ages 12 years from now will be 89 years. Find their ages
Answer by ikleyn(52790)

.

    (Jun + 12) + (July + 12) = 89


    ((x+5)+12) + (x+12)      = 89


          2x                 = 89 - 5 - 2*12 = 60

           x                                 = 60/2 = 30.


ANSWER.  July is 30 years old.  Jun is 35 years old.